OpenStudy (eva12):

1)A satellite in a circular orbit 700 miles above the surface of the Earth. What is the period of the orbit? You may use the following constants: Radius of the Earth: 4000 miles Gravitational Constant: 6.67×10^−11m3/(kg⋅s2) Mass of earth: 5.98×10^24kg Number of Meters in a mile: 1609 Period = seconds help please i tried this question a few times but cant get it right i use w=sqrt(GM)/R^(3/2)

OpenStudy (jack1):

period = time for one complete osciliation si first question: is it geostationary? if so...24 hrs...? so period in seconds is \[\frac {24~hrs}{1} \times \frac {60~mins}{1~hr} \times \frac {60~sec}{1~min} = x~sec\] so 24 x 60 x 60 = number of seconds